A Bibliography of TEXniques Publications

Nelson H. F. Beebe Center for Scientific Computing

University of Utah

Department of Mathematics, 322 INSCC

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

FAX: +1 801 585 1640, +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org, beebe@ieee.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 November 2000 Version 1.18

Abstract

This bibliography summarizes the TEX User Group TEXniques series, special publications about the use of TEX and .

Title word cross-reference

Approximate [Gol88]. Art [Wit88].

Between [Far88]. Books [Kru88].

C [Lev87a]. CapTEX [Sch88]. Case [Jul88]. Choosing [Far88]. Chose [TW88]. Club [Har88]. Come [Chi87]. Command [BB88]. Commercial [Nes87]. Courses [C^{+}88].

Databases [NS88]. Developing [McG87]. Development [Ren88]. Do [SK87].

Documentation [Pol88, Pri88]. Driver [Bee87, McG87]. DVI [Bee87, McG87, Yap88].

Edit [vB88]. Editor [McP88, Mul88, vB88]. Emacs [vB88]. End [Mul88]. Environment [Jul88, Nes87, Ren88]. Experience [Moo88]. Experiences [Saa88].

Family [Bee87]. FASTEX [Mul88]. Finland [Saa88]. First [Sam83]. Formatting [Dye87]. Front [Mul88]. Front-End [Mul88].

Gentle [Doo88]. Getting [Par88]. GloTEX [Aur86]. Government [Pol88]. Grade [Sam83]. Greek [Lev87b]. Guide [Aur87, Aur86, Hut87, McP88].

IBM [Sim87]. IdxTEX [Aur87].

Implementing [Jul88]. Industrial [Sch88]. INRSTEX [Fer86]. Integrated [Ren88]. Introduction [Doo88, Urb90].

Japanese [KS88, SK87, Sai87]. Journal [Thi87]. JTEX [Sai87].

Kennel [Har88].

Language [McP88]. Language-Sensitive [McP88]. Languages [Gol88]. Lawyer [Dye87]. Layout [BN88]. Linguistics [Thi87]. Literate [Lev87a]. Long [Chi87]. LSEDIT [McP88].

Making [Fer86]. Manual [Man88, Wic87]. Maryland [Dye87]. Mathematics [KS88]. Method [Fer86]. ware [KRS90]. Multilingual [Fer88].

NASA [McC88].

Ottoman [AM87].

PC [Mul88]. PiCTEX [Wic87]. Porting [Sim87]. PreTEX [Kru88]. Previewers [Yap88]. Produce [Har88, Pol88]. Producing [Man88, McC88]. Production [Jul88, Moo88, Thi87, Wit88]. Program [Aur87, Aur86]. Programming [Lev87a]. Project [AM87]. Publishing [KS88].

Quick [McP88].

Reference [McP88]. Reports [McC88]. RT [Sim87].

Safely [vB88]. Semitic [Gol88]. Sensitive [McP88]. Sets [Man88]. SGML [Pri88]. Should [SK87]. Single [Man88]. Single-Sourcing [Man88]. Sources [vB88]. Sourcing [Man88]. Standard [Pol88]. Standards [McG87]. Started [Par88]. Strength [Sch88]. Study [Jul88].

Summary [BB88]. Syllabi [C^{+}88]. System [Ren88].

Table [Fer86]. Teaching [Wit88]. Technical [Kru88, McC88]. Text [Dye87, Mul88]. Textbook [KS88, Moo88]. Texts [AM87]. Tips [Par88]. Tools [Kru88]. Trade [TW88]. Typesetter [TW88]. Typesetting [Kru88, Lev87b].

Use [Nes87, Ren88]. User

[Aur87, Aur86, Hut87, Pri88]. Using [Har88, Man88, Pol88, Pri88, vB88].

VAX [McP88].

Way [Chi87]. We’ve [Chi87].

Yearbooks [Har88].

References

Andrews:TQ5-35

[AM87] Walter Andrews and Pierre MacKay. The Ottoman Texts Project. TEXniques, Publications for the TEX community, (5):35–52, 1987.

Aurbach:TQ4-1

[Aur86] R. L. Aurbach. User’s guide to the GloTEX program. TEXniques, Publications for the TEX community, (4):i, 1–10, 1986.

Aurbach:TQ3-1

[Aur87] R. L. Aurbach. User’s guide to the IdxTEX program. TEXniques, Publications for the TEX community, (3):i, 1–14, 1987.

Botway:TQ10-1

[BB88] L. Botway and C. Biemesderfer. LaTEX command summary. TEXniques, Publications for the TEX community, (10):1–14, 1988.

Beebe:TQ5-71

[Bee87] Nelson H. F. Beebe. A TEX DVI driver family. TEXniques, Publications for the TEX community, (5):71– 114, 1987.

Barnhart:TQ7-97

[BN88] Elizabeth Barnhart and David Ness.

Layout for TEX. TEXniques, Publications for the TEX community, (7): 97–116, 1988.

Childs:TQ7-117

[C^{+}88] Bart Childs et al. Syllabi for TEX and courses. TEXniques, Publications for the TEX community, (7): 117–128, 1988.

Carr:TQ5-141

[Car87] Leslie Carr. Of and PostScript. TEXniques, Publications for the TEX community, (5): 141–152, 1987.

Childs:TQ5-1

[Chi87] Bart Childs. We’ve come a long way, and ? TEXniques, Publications for the TEX community, (5):1–4, 1987.

Doob:TQ12-1

[Doo88] Michael Doob. A gentle introduction to TEX. TEXniques, Publications for the TEX community, (12):iv, 1–88, 1988.

Dyer:TQ5-133

[Dye87] Allen R. Dyer, Esquire. Text formatting and the Maryland lawyer. TEXniques, Publications for the TEX community, (5):133–140, 1987.

Farrell:TQ7-155

[Far88] Shawn Farrell. Choosing between TEX and LaTEX. TEXniques, Publications for the TEX community, (7): 155–164, 1988.

Ferguson:TQ2-1

[Fer86] Michael J. Ferguson. Table making —the INRSTEX method. TEXniques, Publications for the TEX community, (2):1–14, 1986.

Ferguson:TQ7-179

[Fer88] Michael J. Ferguson. TEX is multilingual. TEXniques, Publications for the TEX community, (7):179–188, 1988.

Goldberg:TQ7-171

[Gol88] Jacques Goldberg. Approximate TEX for Semitic languages. TEXniques, Publications for the TEX community, (7):171–178, 1988.

Harris:TQ7-83

[Har88] Robert L. Harris. Using TEX to produce kennel club yearbooks. TEXniques, Publications for the TEX community, (7):83–96, 1988.

Huth:TQ8-1

[Hut87] Frances Huth. A users’ guide for TEX. TEXniques, Publications for the TEX community, (8):v, 1–51, 1987.

Jul:TQ7-53

[Jul88] Eric Jul. Implementing TEX in a production environment: A case study. TEXniques, Publications for the TEX community, (7):53–60, 1988.

Knuth:TQ13-1

[KRS90] Donald E. Knuth, Tomas G. Rokicki, and Arthur Samuel. ware. TEXniques, Publications for the TEX community, (13):iv, 101–126, 201– 239, 301–380, 401–441, 1990.

Kruse:TQ7-219

[Kru88] Robert L. Kruse. PreTEX: Tools for typesetting technical books. TEXniques, Publications for the TEX community, (7):219–226, 1988.

Kitagawa:TQ7-165

[KS88] Kazuhiro Kitagawa and Nobuo Saito.

Mathematics textbook publishing with Japanese TEX. TEXniques, Publications for the TEX community, (7): 165–170, 1988.

Levy:TQ5-125

[Lev87a] Silvio Levy. Literate programming in C. TEXniques, Publications for the TEX community, (5):125–130, 1987.

Levy:TQ5-27

[Lev87b] Silvio Levy. Typesetting Greek.

TEXniques, Publications for the TEX community, (5):27–34, 1987.

Mann:TQ7-31

[Man88] Laurie Mann. Producing manual sets using single-sourcing. TEXniques, Publications for the TEX community, (7):31–40, 1988.

McCaskill:TQ7-1

[McC88] Mary McCaskill. Producing NASA technical reports with TEX. TEXniques, Publications for the TEX community, (7):1–10, 1988.

McGaffey:TQ5-69

[McG87] Robert W. McGaffey. Developing TEX DVI driver standards. TEXniques, Publications for the TEX community, (5):69–70, 1987.

McPherson:TQ1-1

[McP88] Kent McPherson. VAX Language-Sensitive Editor (LSEDIT) Quick Reference Guide. TEXniques, Publications for the TEX community, (1): ii, 1–9, 1988.

Mooney:TQ7-69

[Moo88] James D. Mooney. An experience in textbook production. TEXniques, Publications for the TEX community, (7):69–82, 1988.

Muller:TQ7-235

[Mul88] Paul M. Muller. FASTEX: A PC text editor and front-end for TEX. TEXniques, Publications for the TEX community, (7):235–254, 1988.

Ness:TQ5-115

[Nes87] David Ness. The use of TEX in a commercial environment. TEXniques, Publications for the TEX community, (5):115–124, 1987.

Ness:TQ7-25

[NS88] David Ness and James Slagle. TEX and databases. TEXniques, Publications for the TEX community, (7):25– 30, 1988.

Parks:TQ7-129

[Par88] Berkeley Parks. TEX tips for getting started. TEXniques, Publications for the TEX community, (7):129–148, 1988.

Pollari:TQ7-41

[Pol88] Jean Pollari. Using TEX to produce government standard documentation. TEXniques, Publications for the TEX community, (7):41–52, 1988.

Price:TQ7-203

[Pri88] Lynne A. Price. Using SGML and TEX for user documentation. TEXniques, Publications for the TEX community, (7):203–210, 1988.

Renfrow:TQ7-11

[Ren88] J. Tom Renfrow. Use of TEX in an integrated system development environment. TEXniques, Publications for the TEX community, (7):11–24, 1988.

Saarinen:TQ7-189

[Saa88] Kauko Saarinen. Experiences with TEX in Finland. TEXniques, Publications for the TEX community, (7): 189–194, 1988.

Saito:TQ5-57

[Sai87] Yasuki Saito. Japanese TEX: JTEX.

TEXniques, Publications for the TEX community, (5):57–68, 1987.

Samuel:TQ11-1

[Sam83] Arthur Samuel. First grade TEX.

TEXniques, Publications for the TEX community, (11):1–34, 1983.

Schmidt:TQ7-227

[Sch88] Mike Schmidt. CapTEX: Industrial strength TEX. TEXniques, Publications for the TEX community, (7): 227–234, 1988.

Simpson:TQ5-131

[Sim87] Richard Simpson. Porting TEX to the IBM RT. TEXniques, Publications for the TEX community, (5):131–132, 1987.

Saito:TQ5-53

[SK87] Nobuo Saito and Kazuhiro Kitagawa. What should we do for Japanese TEX. TEXniques, Publications for the TEX community, (5):53–56, 1987.

Thiele:TQ5-5

[Thi87] Christina Thiele. TEX, linguistics, and journal production. TEXniques, Publications for the TEX community, (5):5–26, 1987.

Tonkin:TQ7-61

[TW88] Peter Tonkin and Alex Warman. How and why a trade typesetter chose TEX. TEXniques, Publications for the TEX community, (7):61–68, 1988.

Urban:TQ9-1

[Urb90] Michael Urban. An introduction to LaTEX. TEXniques, Publications for the TEX community, (9):iii, 1–56, 1990.

vonBechtolsheim:TQ7-195

[vB88] Stephan von Bechtolsheim. Using the Emacs editor to safely edit TEX sources. TEXniques, Publications for the TEX community, (7):195–202, 1988.

Wichura:TQ6-1

[Wic87] Michael J. Wichura. The PiCTEX manual. TEXniques, Publications for the TEX community, (6):v, 1–85, 1987.

Wittbecker:TQ7-149

[Wit88] Alan Wittbecker. The art of teaching TEX for production. TEXniques, Publications for the TEX community, (7): 149–154, 1988.

Yap:TQ7-211

[Yap88] Ken Yap. DVI previewers. TEXniques, Publications for the TEX community, (7):211–218, 1988.